OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}a_nx^{n-1}\] so now if n=2 would the sum change to \[\sum_{n=2}^{\infty}a_{n+1}x^n\]?

OpenStudy (anonymous):



OpenStudy (anonymous):

so I did it wrong?

OpenStudy (anonymous):

why is it \[a_{n-1}\]?

zepdrix (zepdrix):

Is this for DE, power series? Trying to get the terms to match so you can combine them?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I posted the whole problem earlier, but no one wanted to help me, so I broke it up term by term

zepdrix (zepdrix):

AHHHHH OpenStudy froze again! Lost my message :( gimme a sec to retype... sigh

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

I find it helpful to make a substitution for n, instead of trying to do it all in your head.\[\large \text{Let} \quad k=n-1, \quad \text{then} \quad n=k+1\] \[\large \sum_{n=1}^\infty a_n\cdot x^{n-1} \qquad \rightarrow \qquad \sum_{k=0}^\infty a_{k+1}\cdot x^k\]

zepdrix (zepdrix):

The subscript on A, and the power on X should be pretty straightforward. Do you understand where I plugged those in? Then for the lower limit of our sum, when n = 1, it looks like based on our substitution, \[\large k=1-1\]

zepdrix (zepdrix):

We want our X term to be x^n or x^k, no additional +1 or -1 nonsense. So that is how we determine the substitution. But I'm sure you already knew that, since you were shifting the N value over 1 place.

OpenStudy (anonymous):

give me a min to process this (or two ;P)